Determine the monotony of the following sequences.

a) Un = 2n+3 / 5n-1 n € N

b) Un = √(n+3) - √(n+1) n € N*

c) Un = Σ from r=1 to n (1/ n+r) , n € N


1
Expert's answer
2021-09-29T00:44:29-0400

"\text{We have that a function is monotone increasing if its derivative is positive and }\\
\text{and monotone decreasing if its derivative is negative. Therefore we find the derivative}\\
\text{of each of the given functions}\\
\text{For the first function $U_n= \frac{2n +3}{5n-1}$, its derivative is}\\
\qquad \qquad \qquad \qquad \qquad 2-\frac{3}{5n^2}\\
\text{which is positive for all $n \in \mathbb{N}$, hence the function is monotone increasing}\\
\text{For the second function $U_n = \sqrt{n+3}-\sqrt{n+1}$, its derivative is}\\
\qquad \qquad \qquad \qquad \qquad \frac{1}{2\sqrt{n+3}}-\frac{1}{2\sqrt{n+1}}\\
\text{which is negative for all $n \in \mathbb{N}$, hence the function is monotone decreasing}\\

\text{For the third function $U_n = \sum^n_{r=1}\frac{1}{n+r}$, to check for the monotonicity, we compute}\\\text{ $U_n-U_{n-1}= \frac{1}{n+1}-\frac{1}{2n+1}-\frac{1}{2n+2}$ }\\
=\frac{1}{2n+2}-\frac{1}{2n+1}\\
=\frac{-1}{(2n+2)(2n+1)}<0\\
\text{Since $U_n-U_{n-1}<0 \implies U_n < U_{n-1}$, the fuction is monotone decreasing}"